### v1

# AutoMisty: A Multi-Agent LLM Framework for Automated Code Generation in the Misty Social Robot

Abstract
The social robot’s open API allows users to customize open-domain interactions. However, it remains inaccessible to those without programming experience. In this work, we introduce AutoMisty, the first multi-agent collaboration framework powered by large language models (LLMs), to enable the seamless generation of executable Misty robot code from natural language instructions. AutoMisty incorporates four specialized agent modules to manage task decomposition, assignment, problem-solving, and result synthesis. Each agent incorporates a two-layer optimization mechanism, with self-reflection for iterative refinement and human-in-the-loop for better alignment with user preferences. AutoMisty ensures a transparent reasoning process, allowing users to iteratively refine tasks through natural language feedback for precise execution. To evaluate AutoMisty’s effectiveness, we designed a benchmark task set spanning four levels of complexity and conducted experiments in a real Misty robot environment. Extensive evaluations demonstrate that AutoMisty not only consistently generates high-quality code but also enables precise code control, significantly outperforming direct reasoning with ChatGPT-4o and ChatGPT-o1. All code, optimized APIs, and experimental videos will be publicly released through the webpage: AutoMisty.
I INTRODUCTION
Social robots are becoming increasingly prevalent in homes, workplaces, and educational settings, providing companionship, assistance, and engagement across various applications. Misty, a versatile social robot, is gaining adoption for interactive and assistive roles. For example, it can serve as a mental well-being coach [1], a caregiver for the elderly [2], or a learning companion for children [3]. As these robots evolve to support diverse interactions, customizing their behavior to align with individual preferences remains a challenge. Despite the availability of open APIs and developer-friendly frameworks, personalization often requires advanced programming skills, making it inaccessible to non-technical users [4, 5].
Large Language Models (LLMs) are emerging as a promising solution for enhancing human-robot interaction due to their advanced code generation capabilities. Recent research has explored leveraging LLMs to improve programming systems, reducing the need for extensive coding skills [6, 7]. However, a fully automated, end-to-end system has yet to be realized. At the same time, existing LLM-driven frameworks face significant challenges, including reasoning limitations, hallucinations, and execution reliability issues, leading to frequent errors in robot programming and action execution [8]. Additionally, the absence of robust feedback mechanisms and API integration often results in unreliable task execution, as many LLMs struggle to retain essential contextual information over prolonged interactions [9]. Recent advancements in multi-agent LLM frameworks have attempted to address these issues by incorporating structured reasoning and collaborative planning for more reliable task execution [10]. However, further optimization is still required to ensure adaptability and responsiveness within a fully end-to-end system.
Figure 1: An overview of the workflow comparison: (a) the previous way to customizing human-robot interaction and (b) the AutoMisty framework for enhanced customization. AutoMisty enables autonomous task planning, task assignment, self-reflection, and executable code generation while adapting to user preferences.
To tackle these challenges, we introduce AutoMisty, illustrated in Fig.1, a novel multi-agent LLM framework that enhances human-robot interaction. It enables autonomous task planning, task assignment, self-reflection, and executable code generation, all while adapting to user preferences. Our key contributions are as follows:
  • We present AutoMisty, the first multi-agent system designed for natural language-driven robot programming. It autonomously formulates task plans, performs reasoning, and generates executable code, allowing Misty to execute complex instructions with user-specific adaptability.
  • We design a two-layer reasoning mechanism for better decision-making and execution reliability: (i) self-reflection for iterative improvement and (ii) human-in-the-loop feedback for continuous adaptation. Additionally, we optimize the APIs to enhance interpretability for LLMs.
  • We construct a 28-task benchmark spanning different complexity levels and introduce five evaluation metrics. Extensive experiments and ablation studies showcase the effectiveness and efficiency of AutoMisty.
Figure 2: Overview of the AutoMisty Framework. Given an open-domain task description, the Planner Agent decomposes the task, formulates a plan, and determines the execution order. It then routes subtasks to specialized agents (Action, Touch, or Audiovisual). Leveraging the optimized Misty API, each subtask agent refines details and generates executable code. Each agent incorporates a two-layer optimization mechanism: Layer 1 involves self-reflection with iterative feedback from a designer-critic interaction, while Layer 2 integrates human-in-the-loop feedback for continuous adaptation. Memory ensures consistency across future generations. The bottom illustrates the interaction process for this task.
II RELATED WORKS
Advancements in LLMs and natural language-driven programming have accelerated interactive multi-agent systems for robotics. AutoMisty leverages AI-driven code generation, multi-agent collaboration, and interactive programming to decompose human instructions into executable robotic control.
AI for Code Generation: Recent AI advancements enable automated code synthesis from natural language. State-of-the-art methods enhance performance through planning mechanisms [11] or code-document retrieval [12]. Multi-turn synthesis, as in CodeGen [13], improves iteration but introduces challenges: reliance on predefined test cases limits generalization, documentation retrieval depends on external sources, and multi-turn refinement can compound errors. AutoMisty addresses these issues by integrating real-time user feedback and a multi-agent system, where specialized agents validate intermediate outputs, ensuring robustness, reducing cascading errors, and enhancing adaptability.
LLM Agents for Planning and Reasoning: Large language models (LLMs) serve as powerful tools for structured planning and reasoning. For instance, the self-consistency approach [14] suggests repeatedly generating potential solutions and using majority voting to finalize the answer, while ReAct [15] employs explicit reasoning and grounding to further enhance an LLM’s capacity for problem sovling. Iterative optimization techniques—such as self-explanation [16], language reinforcement [17], tree of thought [18], and chain of thought [19]—build upon past decisions to refine learning outcomes. Concurrently, state-based modeling [20], curriculum-driven learning [21], and generative agents [22] promote adaptive reasoning, collectively driving progress in task planning. Drawing inspiration from these methods, Automisty integrates human-in-the-loop validation [20] to ensure the reliability of outputs and mitigate error propagation. Furthermore, it introduces a Critic module for automated reflection [17] and continual optimization, streamlining collaborative workflows.
LLMs for Robotic Control: LLMs have demonstrated significant potential in robotic control by leveraging vision-language-action models and interactive programming. Key advancements include repurposing code-generation models for robot policy synthesis [23, 24, 25, 26], extracting actionable knowledge without retraining [27, 28, 29, 30, 31], integrating foundation models for hierarchical problem-solving [32], and incorporating embodied language models (ELMs) with continuous sensor data [33]. However, these approaches either suffer from hallucination errors or require extensive pretraining. Automisty addresses these challenges through interactive programming, where a compiler verifies code correctness, and human observation provides real-world feedback on Misty’s performance, effectively eliminating hallucination issues. Additionally, the system employs in-context learning, circumventing the need for model retraining.
Framework for Multi-Agent LLM Systems: The Multi-Agent LLM Framework abstracts LLM models into distinct agents, each assigned specialized roles, making it easier to design systems with multiple collaborative agents. Existing frameworks facilitate multi-agent interactions [34, 35], simulate software development workflows [36, 37], or generate autonomous agents [38]. Automisty is built upon the Autogen framework as its foundational implementation [34].
III PROBLEM FORMULATION
Given a natural language task $\mathcal{I}$, AutoMisty aims to generate executable MistyCode code $C$ to accomplish the given task. Initially, the system performs semantic analysis and requirement delineation on $\mathcal{I}$, then breaks it down into several subtasks $T=T_{1},T_{2},...,T_{k}$. If a subtask is too complex, it can be further subdivided until each subtask is sufficiently atomic to execute. The system then leverages a set of $N$ large language model Agents, denoted by $A$, to handle each subtask. Each Agent in this set specializes in a different skill set to handle a specific subtask. A unique assignment function then maps each subtask $T_{j}$ to an Agent that meets its skill requirements. This is defined as:
$\phi:\{\,T_{1},\dots,T_{k}\}\;\to\;A$
If the output of Agent $A_{j}$ is required as input for Agent $A_{i}$, a directed edge $(A_{j}\to A_{i})$ is added to a directed acyclic graph $G$, which represents this multi-agent collaboration. As $G$ is acyclic, a topological sort produces the execution sequence:
$\pi(G)\;=\;\langle A_{i_{1}},\,A_{i_{2}},\,\dots,\,A_{i_{m}}\rangle,$
ensuring that for any Agent in the sequence, all of its upstream dependencies have finished before it executes. For the $k$-th Agent in the execution order, $A_{i_{k}}$, its input condition, $\Delta(A_{i_{k}})$ comprises all the code or contextual information $C$ generated by its upstream Agents. We then define the code generation function
$\Gamma(A_{i_{k}}\,\mid\,\Delta(A_{i_{k}}))\;=\;C_{i_{k}},$
where $C_{i_{k}}$ is the code that $A_{i_{k}}$ produces upon completing its subtask. Once the last agent in the sequence is reached, its output becomes the final MistyCode $C$ that completes the task $\mathcal{I}$.
TABLE I: List of tasks under each category
Elementary Task (1) Hand Raise Direction Change, (2) Hand Raise Count Change, (3) Hand Raise Speed Change, (4) Head Turn Direction Change, (5) Head Turn Count Change, (6) Head Turn Speed Change, (7) Expression Transformation Test, (8) Speech Speed Change Test, (9) LED Color Change Test
Simple Task (10) Storytelling, (11) Timed Reminder, (12) Emotional Comfort, (13) Poetry Composition and Performance, (14) Key Word Triggered Video Recording, (15) Rock-style Dance Performance, (16) Key Word Triggered Photography
Compound Task (17) Reacting to Head Touch with Pikachu Noise, (18) Expressing Different Emotions Based on Touch Location, (19) Recognizing Colors and Displaying with LED, (20) Waving in Response to a Greeting
Complex Task (21) Recognizing and Reading Signs, (22) Simulating a Child’s Persona in Conversation, (24) Math Tutoring, (24) Intelligent Storytelling, (25) Real-Time Translation, (26) Taking a Photo and Performing Analysis, (27) Countdown Before Taking a Photo and Performing an Action (28) Playing a Game Like Concentration
IV METHODOLOGY
This section outlines the AutoMisty methodology. We first introduce the optimized Misty APIs and their assignments, which are essential for enhancing LLM comprehension and helping agents focus on their designated tasks. Next, we present the two-layer optimization mechanism within each agent, ensuring reliable and well-structured code. Finally, we describe the collaboration among the four agents utilized in our proposed method and detail their specific functions. Fig. 2 illustrates this approach.
IV-A Optimized Misty APIs
We observe several issues with Misty’s APIs, including unclear function descriptions, missing parameter definitions, and incomplete interfaces, which often lead to errors and hallucinations during execution. To ensure high-quality and efficient code generation, an optimized API is essential.
Inspired by Song et al. [39], we find that detailed code annotations greatly enhance LLM accuracy in code generation and execution. We first analyze Misty’s capabilities, identifying three core task categories: action, touch, and audiovisual processing. We then refine and restructure the APIs, improving audio-video data flow and integrating OpenAI’s Whisper model for real-time translation, enabling Misty to “see” and “hear.” Additionally, we encapsulate LLM logic, equipping Misty with basic reasoning and understanding. This process resulted in 136 optimized APIs, for which we provide comprehensive documentation on input/output formats, parameter semantics, application scenarios, and complex functionalities, helping to reduce errors and enhance reliability. Finally, these APIs are assigned to specific subtask agents to improve focus and reduce the likelihood of hallucinations.
IV-B Two-layer Optimization Mechanism
AutoMisty employs a Two-Layer Optimization Mechanism within each agent to enhance reasoning, decision-making, and task execution. When an agent is triggered, it activates a Drafter, which listens to Layer 1’s dialog. Layer 1 utilizes a Self-Reflective Feedback Mechanism, where the Designer iteratively refines the draft using In-Context Learning (ICL), while the Critic evaluates its accuracy and feasibility. If revisions are needed, feedback loops between these two agents until the Critic approves the optimal solution. Once this phase is finalized, the Drafter initiates Layer 2, a human-in-the-loop process where the user provides direct feedback on the current results. If any aspects are deemed unsatisfactory, the Drafter reactivates Layer 1, iteratively refining the output until the user grants final approval. Once the user is satisfied, the agent concludes the current subtask.
The optimization mechanism includes a memory module that enhances the agent’s teachability, allowing it to learn from past successes. When a user saves a successful task, the system refines future code generation while preserving user preferences, improving efficiency and consistency.
IV-C Multi-Agent Collaboration Framework
AutoMisty’s multi-agent collaboration framework consists of four key agents: the Planner Agent, Action Agent, Touch Agent, and Audiovisual Agent, each contributing to different aspects of task execution while ensuring iterative validation and refinement.
The Planner Agent serves as the central coordinator, responsible for analyzing high-level instructions, breaking down tasks into manageable subtasks, and assigning them to the appropriate agents. It generates an action plan with an optimized execution order and ensures coherence before proceeding. It also seeks human validation to align execution with user intent. Once the final plan is generated, the Planner routes it to subtask agents in the designated execution order. The Action, Touch, and Audiovisual Agents execute specialized tasks within their respective domains. Each agent engages in fine-grained reasoning, analyzing relevant APIs, and generating executable code, optionally referencing prior stored memory.
In addition, each agent is designed to isolate their reasoning from each other. This approach allows agents to operate collectively yet independently, ensuring that each agent focuses only on its designated subtask while relying on structured interactions to exchange relevant information. This isolation prevents any agent from being overwhelmed with excessive context, thus improving efficiency and decision-making accuracy.
IV-D System Verification and User Proxy
Once the MistyCode has been generated, AutoMisty verifies the code in a local environment before forwarding it to Misty for execution. If compilation or runtime errors occur, the system captures exception details and feeds them back to the Designer, enabling debugging and refinement within an updated context that includes error messages and partial execution results. This execution-feedback loop repeats until the code runs successfully in the local environment. Once the code passes verification, Misty executes the corresponding instructions. Inspired by AutoGen [34], we utilize the User Proxy to proxy user input to the system, allowing the user to observe the robot’s actual behavior and propose further refinements as needed.
. Tasks ChatGPT_4o ChatGPT_o1 AutoMisty (Ours) TC NUI CE US TC NUI CE US TC NUI CE US UPI TCI UPI TCI UPI TCI Elementary Tasks 1 - 4 Success 0 0 10 10 Success 0 0 10 10 Success 0 0 10 10 5 Success 1 0 10 10 Success 0 1 10 10 Success 1 0 10 10 6 Fail - - - - Success 0 0 10 10 Success 0 0 10 10 7 Success 2 1 10 10 Success 0 1 10 10 Success 1 0 10 10 8 Success 0 0 10 10 Success 0 0 10 10 Success 0 0 10 10 9 Success 0 0 10 10 Success 0 1 10 10 Success 1 0 10 10 Simple Tasks 10 Success 1 0 10 10 Success 1 0 10 10 Success 1 0 10 10 11 Success 0 0 10 10 Success 0 0 10 10 Success 0 0 10 10 12 Success 1 0 10 10 Success 0 0 10 10 Success 1 0 10 10 13 Success 0 0 10 10 Success 0 0 10 10 Success 0 0 10 10 14 Fail - - - - Fail - - - - Success 0 0 10 10 15 Success 0 0 10 10 Success 0 0 10 10 Success 0 0 10 10 16 Fail - - - - Success 0 0 10 10 Success 0 0 10 10 Compound Tasks 17 Fail - - - - Success 1 0 10 10 Success 0 1 10 10 18 Fail - - - - Fail - - - - Success 1 0 10 10 19 Fail - - - - Success 1 0 10 10 Success 1 0 9 10 20 Fail - - - - Success 1 0 10 10 Success 1 0 9 10 Complex Tasks 21 Fail - - - - Success 0 0 10 10 Success 0 1 9 10 22 Success 0 0 10 10 Success 0 0 10 10 Success 0 0 9 10 23 Fail - - - - Success 0 0 10 10 Success 0 0 9 10 24 Success 0 0 9 10 Success 0 0 10 10 Success 0 0 10 10 25 Fail - - - - Success 0 0 10 10 Success 0 0 9 10 26 Fail - - - - Fail - - - - Success 0 0 9 10 27 Fail - - - - Fail - - - - Success 1 2 10 10 28 Fail - - - - Success 6 4 10 8 Success 3 2 9 10
TABLE II: Performance Comparison of ChatGPT-4o, ChatGPT-o1, and AutoMisty Across Different Categories of Tasks
V EXPERIMENTS
V-A Benchmark Dataset
To evaluate AutoMisty’s performance and compare it with baseline methods, we created a benchmark dataset for natural language-based task code generation and execution with the Misty robot. We invited 10 people (3 engineers and 7 non-engineers) to suggest tasks for Misty, merging overlapping ones. The dataset includes tasks of varying complexity, grouped into four categories:
  • Elementary Tasks include basic commands such as raising the left or right hand, adjusting head angles, and controlling movement speed. The primary goal is to test the user’s ability to precisely control the robot through conversation.
  • Simple Tasks are tasks that can be completed using a single agent with minimal creativity. The solution is straightforward, and execution is direct.
  • Compound Tasks require multiple agents to collaborate in completing the task. The instructions are clear, and the interaction between agents follows a well-defined structure.
  • Complex Tasks are highly abstract and vague instructions that require multiple agents working in tandem. The system must autonomously design LLM prompts to accomplish the task. Solving the task demands significant creativity, potentially involving external tool integration or even the design of new tools.
The dataset consists of 28 tasks in total: 9 elementary tasks, 7 simple tasks, 4 compound tasks, and 8 complex tasks. Table I provides a list of tasks under each category.
TABLE III: Performance Comparison of AutoMisty With vs. Without Self-Reflection
Tasks AutoMisty (W/O Self-Reflection) AutoMisty (W Self-Reflection)
TC NUI CE US TC NUI CE US
UPI TCI UPI TCI
Elementary Tasks
1 - 9 Success 0 0 10 10 Success 0 0 10 10
Simple Tasks
10 Success 1 0 10 10 Success 1 0 10 10
11 Success 1 0 10 10 Success 1 0 10 10
12 Success 0 0 10 10 Success 0 0 10 10
13 Success 0 2 10 10 Success 0 0 10 10
14 Success 0 0 10 10 Success 0 0 10 10
15 Success 0 0 10 10 Success 0 0 10 10
16 Success 0 0 10 10 Success 0 0 10 10
Compound Tasks
17 Success 0 0 10 6 Success 0 1 10 5
18 Success 0 0 10 10 Success 0 1 10 10
19 Success 1 1 10 10 Success 1 0 9 10
20 Success 0 0 10 10 Success 1 0 9 10
Complex Tasks
21 Fail - - - - Success 0 1 9 10
22 Fail - - - - Success 0 0 9 10
23 Success 0 0 10 10 Success 0 0 9 10
24 Success 0 0 10 10 Success 0 0 9 10
25 Success 0 0 10 10 Success 0 0 9 10
26 Success 0 0 10 10 Success 0 0 9 10
27 Success 0 0 10 10 Success 1 2 10 10
28 Fail - - - - Success 3 2 9 10
V-B Evaluation Metrics
We evaluate performance using the following metrics: Task Completion (TC) status, Number of User Interactions (NUI), Code Efficiency (CE) score, and User Satisfaction (US) rating. Within NUI, we further assess User Preference Interactions (UPI) and Technical Correctness Interactions (TCI).
  • TC indicates whether the task is completed within 20 interactions.
  • NUI represents the total number of user interactions during the process.
  • UPI counts interactions where the user participated in an interaction due to personal preference.
  • TCI tracks interactions where the user provides suggestions related to coding correctness or implementation logic.
  • CE evaluates the efficiency of the generated code. If a portion of the code can be removed without affecting correctness, a penalty of -1 is applied. The maximum possible score is 10.
  • US measures user satisfaction with the generated code. If an interaction’s instructions are misinterpreted or not followed, a penalty of -1 is applied. The maximum possible score is 10.
VI RESULTS AND DISCUSSION
This section presents experimental results with AutoMisty and baseline methods on unseen dataset tasks, as summarized in Table II.
We observe that for elementary and simple tasks, all models achieve a high success rate with minimal user interaction. However, as task complexity increases, we find that AutoMisty and ChatGPT-o1 begin to outperform ChatGPT-4o. The most significant differences emerge in complex tasks, where AutoMisty shows the highest robustness and adaptability. ChatGPT-4o failed most complex tasks (4 out of 8) while ChatGPT-o1 failed a few (2 out of 8). Although AutoMisty’s CE is a little lower than that of ChatGPT-o1, it was able to complete all tasks successfully. This demonstrates AutoMisty’s capability to generate MistyCode across all task complexities.
Performance Variation Analysis: The non-deterministic behavior of LLMs leads to inherent variability in their outputs [40]. To analyze this, we performed five independent trials, selecting the tasks that both baseline models, ChatGPT-4o and ChatGPT-o1, failed to complete. Table IV reports the mean and standard deviation of the results obtained across these trials for our approach. We observe that our proposed method consistently produces similar outcomes across all tasks.
TABLE IV: AutoMisty Performance Variation Analysis
Task TC UPI TCI CE US
14 1.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 10.00$\pm$0.00 10.00$\pm$0.00
18 1.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 10.00$\pm$0.00 10.00$\pm$0.00
26 1.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 10.00$\pm$0.00 10.00$\pm$0.00
27 1.00$\pm$0.00 0.00$\pm$0.00 1.00$\pm$1.00 10.00$\pm$0.00 10.00$\pm$0.00
TABLE V: Teachability Assessment
Emotion Saved in Memory User Input Retrieval Result (Teachability Obseved)
Happiness and Excitement Exhilaration Happiness and Excitement (correct)
Ecstasy Happiness and Excitement (correct)
Jubilation Peaceful Happiness and Satisfaction (incorrect)
Rage Fury Rage (correct)
Wrath Rage (correct)
Outrage Rage (correct)
Grief Sadness Grief (correct)
Mourning Grief (correct)
Heartache Grief (correct)
Peaceful Happiness and Satisfaction Tranquil Bliss Peaceful Happiness and Satisfaction (correct)
Quiet Euphoria Happiness and Excitement (incorrect)
Placid Joy Peaceful Happiness and Satisfaction (correct)
Impact of the Self-Reflective Feedback Mechanism: To analyze the impact of the Self-Reflective Feedback Mechanism on AutoMisty’s performance, we conduct experiments across all four task complexities without self-reflection. These results are highlighted in Table III. We find that both versions of AutoMisty performed consistently well on both elementary and simple tasks, achieving perfect task completion rates with minimal or no user interactions. In contrast, AutoMisty without the mechanism failed on multiple occasions (Tasks 21, 22, and 28), while the Self-Reflective version was able to complete all tasks. Although Self-Reflection led to a slight increase in the number of interactions (NUI) in complex tasks, it consistently resulted in higher task success rates, demonstrating its overall effectiveness.
Teachability Assessment: Teachability refers to an agent’s ability to learn from past successes. To evaluate this capability in our system, we designed an experiment based on user-specific emotional preferences. Since users perceive emotions uniquely, we stored four well-recognized emotion types in the memory. The user then inputs synonymous emotions, and we assess whether the system correctly retrieves the corresponding anchor emotion.Table V summarizes the results. Although AutoMisty misclassifies some emotions, it accurately retrieves well-defined ones like ”Rage” and ”Grief,” indicating its teachability.
VII CONCLUSION
AutoMisty is the first system to break the technical bottleneck, enabling non-programmers to generate executable code directly through conversational instructions. Our design significantly enhances code control, accuracy, and consistency, achieving a 100% task success rate across all complexity levels, whereas ChatGPT-o1 achieves only 75% on complex tasks. While we cover 28 tasks, our system demonstrates strong generalization capabilities for broader open-domain tasks. By leveraging context optimization strategies, such as API semantic anchoring, the entire framework operates solely through in-context learning, enabling low-cost and efficient migration to other API-driven social robotics. In the future, we will expand our work to include collaboration and interaction between multiple social robots and humans.